# Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21

### v2

@@ Abstract
Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials. Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids and present the results for spheres. We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy. Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
@@
and their unique physical properties. A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4]. Furthermore, by adjusting their geometry, it is often possible to tailor their response at will. In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11]. While isotropic nanostructures have widespread use [12], anisotropic nanostructures are currently at the forefront of photonics research [13, 14].
@@
In addition to this class of resonators, a variety of naturally occurring and existing systems are, in fact, anisotropic particles, ranging from liquid crystal droplets to ferromagnetic particles and ice grains [16, 17, 18, 19, 20, 21]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules. [22, 23, 24, 25].
@@
Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32]. Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36]. Recently, the emergence of anisotropic materials in optics [13, 37, 38, 14, 39] has motivated the investigation of the response of anisotropic inclusions, with works on isotropic-mode-based scattering analysis from a uniaxial sphere, semianalytical and numerical analyses of a uniaxial spheroid, and eigenmode analysis of a biaxial slab [40, 4, 41, 42, 43]. More recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [44].
However, while the dipole response of anisotropic particles is well understood [45], the derivation of higher-order modes and eigenpermittivities remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
@@
Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivity relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source to allow several degrees of freedom of the permittivities in the different axes. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time varying function of the excitation source. In our analysis we show that anisotropic particles exhibit novel radiation patterns. In addition, we identify axial-eigenpermittivity sum
rules, resulting in resonance splitting in anisotropic particles. We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support. We apply our theory to spherical uniaxial and biaxial particles composed of hexagonal boron nitride (hBN) and molybdenum trioxide MoO3) phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 52, 53, 54].
@@
the boundary conditions at the particle interface. In addition to the challenge that the symmetry of
@@
As a result, the widely-employed spherical harmonics cannot be eigenstates
@@
the high-order spherical harmonics are suitable only for isotropic
medium [2, 35], and there
@@
that the anisotropic modes are composed of specific superpositions of isotropic modes.
Based on these considerations, we propose the following first second-order anisotropic sphere modes
where are the isotropic-sphere eigenstates [1, 2]. Inside the sphere
we write
where in order to normalize the modes one has to multiply them by
Imposing Laplace’s equation we obtain
which corresponds to a uniaxial medium inside the sphere.
is continuous at the interface by construction
and we impose continuity of We first calculate inside
the sphere by projecting
where we have used Note that the dot
product with retrieved that were removed in the differentiation,
which gives a factor of 2 as in the isotropic case, only that there
is a freedom of choice of
We then calculate outside the sphere
By equating them we obtain
Note that the resonance condition depends on which can be utilized to sense the environment in which the particle is situated.
Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere. (a) A uniaxial sphere and the permittivity of bulk hBN. (b) The scattering spectrum and of the dipole modes on an hBN sphere surface. Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively. The crystal axes [100],[010],[001] are directed along the axes, respectively. (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material. (d) Scattering spectrum and of the third-order mode, where is excited at two s. In addition, there are geometric-anisotropic shifts compared to the bulk for all modes. Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
We now suggest the second second-order mode for an anisotropic sphere
and write the potential and field inside the sphere
where to normalize the modes one has to multiply them by Laplace’s equation is satisfied for this mode
We calculate at the boundaries
By equating these expressions we get
which for reads
Since the choice of system of coordinates is arbitrary, the derived modes
and eigenpermittivity relations should hold for any two axes. The
@@
this time for a general biaxial medium. Thus, for uniaxial and biaxial spheres four and three second-order modes can be excited, respectively. In addition, one can see that there is a degeneracy
for since both anisotropic
modes are excited. Coherently adding these modes gives for the angular part, where we assumed equal weights of the modes. Moreover, there is an additional degeneracy for a uniaxial sphere arising from the sum rule, and coherently adding the corresponding modes for gives Finally, for a resonance at a given axis, the sum rule is typically approached twice with the two other axes, leading to resonance splitting.
In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions and their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial spheres, which leads to three superimposed modes.
@@
We now proceed to the third-order mode of an anisotropic sphere. We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition. In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the second-order modes, which also display a degeneracy.
@@
We now consider the second-order modes of an anisotropic ellipsoid in analogy to the anisotropic sphere.
@@
We verified the theory for anisotropic ellipsoids with near-field spectroscopy of two nanoellipsoids composed of in Ref. [44] and obtained excellent agreement for both the modes and the spectra. In that work the near-field imaging tip scattered the field component in the axial direction to the far field, which was then measured.
@@
the axial permittivities. We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of the second-order modes of anisotropic particles. Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.

### v3

@@ Abstract
Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials. Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids, and present the results for spheres. We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy. Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
@@
and their unique physical properties. A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4]. Furthermore, by adjusting their geometry, it is often possible to tailor their response at will. In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11]. While isotropic nanostructures have widespread use [12], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [13, 14].
@@
In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [16, 17, 18, 19, 20, 21]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules. [22, 23, 24, 25].
@@
Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32]. Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Recently, the emergence of anisotropic materials in optics [13, 44, 45, 14, 46] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [47, 4, 48, 49, 50, 51, 51, 52, 53, 54].
More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3) [55, 14, 56, 20]. Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
However, while the dipole response of anisotropic particles is well understood [42], the derivation of higher-order modes and eigenpermittivity relations remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
@@
Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [1, 2], to allow several degrees of freedom of the permittivities in the different axes. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source. Note that whereas previous (semianalytical) works on particles with cartesian anisotropy focused on dielectric materials with nonresonant excitations, we consider the resonant modes with at least one [52, 51]. Our analysis shows that anisotropic particles exhibit novel radiation patterns. While we consider the localized modes, in practice the particles have size and radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns [63]. In addition, we identify axial-eigenpermittivity sum
rules, resulting in resonance splitting in anisotropic particles. We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support. We apply our theory to spherical uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 64, 65, 66].
@@
the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system. In addition to the challenge that the symmetry of
@@
As a result, the widely-employed spherical harmonics are in general not eigenstates
@@
the high-order spherical harmonics are valid for isotropic
medium [2, 35],
and there
@@
that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require . Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
It can be seen that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [1, 2] for the equal permittivity axes and new eigenpermittivity relations, respectively. However, our modes in both cases are directional as opposed to the isotropic sphere modes.
Note that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere. (a) A uniaxial sphere and the permittivity of bulk hBN. (b) The scattering spectrum and of the dipole modes on an hBN sphere surface. Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively. The crystal axes [100],[010],[001] are directed alfong the axes, respectively. (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material. (d) Scattering spectrum and of the third-order mode, where is excited at two s. In addition, there are geometric-anisotropic shifts compared to the bulk for all modes. Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of
Thus, we suggest the second-order biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
The
@@
this time for a general biaxial medium.
In addition, one can see that there is a degeneracy
for uniaxial particles with since both are excited. Note that one could select the basis for the uniaxial modes. In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
@@
We now proceed to the third-order biaxial mode of an anisotropic sphere. We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition. In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the second-order modes, which also display a degeneracy.
@@
We now consider the second-order modes for anisotropic ellipsoids in analogy to the anisotropic sphere.
@@
We verified our theory for two biaxial ellipsoids composed of with hyperspectral midIR near-field imaging in Ref. [56] and obtained excellent agreement for both the modes and the spectra. In that work the near-field imaging tip scattered the field component in the axial direction to the far field, which was then measured.
@@
the axial permittivities. We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of anisotropic particles. Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.